Circular No. 7934

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 57P/DU TOIT-NEUJMIN-DELPORTE A faint object reported as a possible NEO candidate yesterday by S.

Pravdo, Jet Propulsion Laboratory, on behalf of NEAT and described as appearing cometary was recognized at the Minor Planet Center as being on the line of variation and to have the motion of comet 57P, but to be displaced from that brighter comet (which was not included in the dataset) by more than 0^{o}.2 on the 1.2-m Palomar Schmidt frames. NEAT had reported the brighter comet on July 1 (see MPEC 2002-N14), but a check with Pravdo yesterday did not reveal anything obvious at the July 1 expected position of the new object. Following a request by the Minor Planet Center, M. Tichý and J. Tichá (Kleť, KLENOT 1.06-m reflector) located the new object last night, describing it as diffuse and having a coma of diameter 8′′–10′′; as did G. Masi and F. Mallia (Campo Catino, 0.8-m f/8 reflector), whose coaddition of five images (for a total integration of 20 min) showed a well-defined 12′′ coma and a delicate northeast-southwest elongation. The object was again included in today’s NEAT NEO-candidate report but with no remark about its appearance.

2002 UT α_{2000} δ_{2000} m_{1} Observer
July 12.20809 20^{h}03^{m}51.^{s}51 −13^{o}43′09′′.8 20.0 Pravdo
12.26994 20 03 50.46 −13 43 12.7 19.2  _{”}
12.33544 20 03 49.31 −13 43 15.7 19.2  _{”}
12.91647 20 03 41.49 −13 43 41.5 19.0 Tichý
12.92166 20 03 41.42 −13 43 43.3  _{”}
12.94456 20 03 40.83 −13 43 45.3  _{”}
12.96377 20 03 40.54 −13 43 45.1  _{”}
13.00443 20 03 39.96 −13 43 44.6 18.7 Masi
13.01694 20 03 39.72 −13 43 46.1 18.7  _{”}
13.03156 20 03 39.52 −13 43 46.8 18.7  _{”}
13.03955 20 03 39.35 −13 43 47.4 19.0  _{”}
13.22201 20 03 37.46 −13 43 52.9 19.6 Pravdo
13.24230 20 03 37.11 −13 43 55.3 20.4  _{”}
13.26421 20 03 36.73 −13 43 56.5 19.6  _{”}

The above positions are fully consistent with the orbital elements for comet 57P on MPC 45964 (ephemeris on MPC 44939) with T for this “component B” delayed by 0.19 day (i.e., to 2002 July 31.37 TT). It should be noted that comet 57P was anomalously bright at its unfavorable 1996 return (T = 1996 Mar. 5.7; m_{1} = 13.3 at the first postperihelic observation on July 24.8—see MPC 27482—some 5 mag brighter than expected).

2002 July 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Brian G. Marsden